OpenStudy (anonymous):

Which of the following trinomials can be factored? a. 3b2+4b+1 b.3b2+6b+14 c. 3b2+8b+1 d. 3b2+30b+2 Please show the work on how you got it factored. Thank you

OpenStudy (amistre64):

any of them can be run thru the quadratic formula; what defines a factoring?

OpenStudy (amistre64):

3.2 = 6; and i cant see and factors of 6 that add up to 30

OpenStudy (amistre64):

or 8

OpenStudy (amistre64):

a looks good for it tho

OpenStudy (amistre64):

and b maybe

OpenStudy (anonymous):

yeah a is correct but can u show me how u factor it. Thanks tips

OpenStudy (amistre64):

3.1 = 3 3+1 = 4 (x+3)(x+1)

OpenStudy (amistre64):

forgot to divide out the 3 again

OpenStudy (amistre64):

(x+3/3)(x+1/3) (x+1)(3x+1)

OpenStudy (amistre64):

thats better

OpenStudy (amistre64):

(x /3)(x /3) 3b2+6b+14 3(14) = 42 i dont think factors of 42 will add to get 6

OpenStudy (amistre64):

5.1=5 4.2=8 3.3=9 none of which are 42

OpenStudy (amistre64):

so i think a is the only factorable one

OpenStudy (anonymous):

so the answer would be (x+1)(3x+1) then?

OpenStudy (amistre64):

yes, for "a"